# Enhancing Clinical Trial Patient Matching through Knowledge Augmentation with Multi-Agents

### v1

@@
Abstract
Matching patients effectively and efficiently for clinical trials is a significant challenge due to the complexity and variability of patient profiles and trial criteria. This paper presents a novel framework, Multi-Agents for Knowledge Augmentation (MAKA), designed to enhance patient-trial matching by dynamically supplementing matching prompts with external, domain-specific knowledge. The MAKA architecture consists of five key components: a knowledge probing agent that detects gaps in domain knowledge, a navigation agent that manages interactions among multiple specialized knowledge augmentation agents, a knowledge augmentation agent that incorporates relevant information into patient-trial matching prompts, a supervision agent aligning the outputs from other agents with the instructions and a matching agent making the final selection decision. This approach enhances the accuracy and contextual richness of patient matching, addresses inherent knowledge gaps in both trail criteria and large language models (LLMs), and improves the alignment between patient characteristics and the criteria.
1 Introduction
Patient matching for clinical trials refers to the process of identifying and enrolling participants whose health profiles align with the specific eligibility criteria of a given clinical study. Efficient patient matching is crucial to accelerating the drug development process and reducing the overall cost of clinical trials. It provides an opportunity for participants to access experimental treatments that could potentially improve their health outcomes and, in some cases, transform their quality of life.
@@ 1 Introduction
The emergence of of Large Language Models (LLMs) offers a promising opportunity to improve patient matching. LLMs, such as GPT-4 Achiam et al. (2023), usually possess a large text window and advanced natural language understanding capabilities (Karanikolas et al., 2023), which enable them process complex trail criteria and patient data effectively. However, their general knowledge may not always be sufficient for knowledge-intensive tasks like patient matching in clinical trials, where domain-specific expertise is required. The knowledge gap inherent in LLMs and misssing information in the pre-defined criteria can limit their performance, highlighting the need for LLM augmentation of specialized medical knowledge (Shuster et al., 2021). While research indicates that LLMs do encode medical knowledge in a parameterized form (Singhal et al., 2023), this information can be outdated or misaligned with the special needs of certain clinical trials.
To address these limitations, we present a new framework called: Multi-Agents for Knowledge Augmentation (MAKA). MAKA is a novel architecture designed to enhance the patient matching process by dynamically integrating domain-specific knowledge aligned with the unique context of clinical trials. It consists of five main components: the Knowledge Probing Agent, Navigation Agent, Knowledge Augmentation Agent, Supervision Agent and Matching Agent. By leveraging this multi-agent framework, MAKA effectively identifies and fills knowledge gaps, enhancing the LLM’s ability to understand complex eligibility criteria and patient profiles. This leads to an average improvement of 3% on accuracy and 6% on F1-score across various patient selection criteria, with up to a 10% boost in more knowledge-intensive cases. Overall, our framework shows significant potentials for achieving more accurate patient matching.
2 Related Work
2.1 Clinical Trail Patient Matching
@@
Traditional patient matching systems rely on query-based structured databases, such as Criteria2Query Yuan et al. (2019), or handcrafted rule-based approaches like RBC Oleynik et al. (2019). However, these methods struggle to handle unstructured free-text data and are highly sensitive to variations in formatting and verbal expressions. Unfortunately, a significant portion of EHRs consists of free-text, including clinical notes, surgical records, radiology reports, pathology reports and so on Sun et al. (2018). Thus, the effectiveness and scalability of traditional methods are significantly limited, making them unsuitable for future patient matching tasks.
@@
Recent advancements in LLMs have shown remarkable capabilities in directly processing free-text clinical trial criteria, offering scalable solutions without the need for extensive rule-based preprocessing. Many Studies have applied these models to patient-centric trial matching tasks using real-world data. For instance, Nievas et al. (2024) fine-tuned the open-source Llama 2 Touvron et al. (2023) to automate patient matching. However, fine-tuning LLMs is computationally expensive, prompting alternative approaches that leverage frozen models. Jin et al. (2023) introduced TrialGPT, a patient-centric matching framework that assesses patient eligibility on a criterion-by-criterion basis, filtering out ineligible trials and ranking potentially eligible ones. Their results indicated a strong correlation between TrialGPT’s predictions and expert clinician judgments, achieving accuracy at the criterion level similar to human experts. Similarly, Gupta et al. (2024) proposed OncoLLM, an end-to-end pipeline that converts trial criteria into conversational prompts and uses a Chain-of-Thought (CoT) Wei et al. (2022) reasoning strategy. OncoLLM outperforms GPT-3.5 while underperforms GPT-4, despite its smaller model size.
@@
In contrast, other studies have focused on trial-centric patient matching, typically using public deidentified datasets due to HIPAA-related data constraints. Wornow et al. (2024) developed a LLM-based system with a two-stage retrieval pipeline to identify suitable patients for trials using a zero-shot prompt approach. It reduces token processing costs while outperforming existing state-of-the-art models on the n2c2 2018 cohort selection benchmark Stubbs et al. (2019). Likewise, Beattie et al. (2024) applied a CoT strategy in a Retrieval-Augmented Generation (RAG) Lewis et al. (2020) system using GPT-3.5 Turbo and GPT-4, demonstrating results comparable to those of Wornow et al. (2024).
@@
These works collectively highlight the potential of LLMs in clinical trial matching. However, existing approaches directly employ LLMs without verifying whether the trial criteria are sufficiently understood by the model Webson and Pavlick (2021) or whether there are gaps in the model’s knowledge Lewis et al. (2020). Our work fits within the trial-centric patient matching literature and makes a novel contribution by automatically identifying and addressing these gaps, thereby enhancing the accuracy patient matching.
2.2 LLM-based Multi-Agents (LLM-MA)
@@
LLM-MA have seen rapid development across various domains, driven by the promising potential of LLMs as autonomous agents. Guo et al. (2024) provide a comprehensive overview of recent advancements in LLM-MA applications, categorizing them into two primary areas: Problem-Solving and World Simulation. Problem-solving LLM-MA systems explore the capabilities of agents with specialized expertise, enabling them to outperform single-agent LLM systems through collaborative efforts. For instance, Wang et al. (2024) proposed Mixture-of-Agents, a layered multi-agent architecture that achieves state-of-the-art performance in benchmarks like AlpacaEval 2.0 Dubois et al. (2024). Similarly, Hong et al. (2023) introduced MetaGPT, a meta-programming framework that integrates human-like task decomposition and workflows, outperforming GPT-4 in various code generation tasks.
In the healthcare domain, the use of LLM-MA systems is emerging. TRIAGEAGENT Lu et al. (2024) utilizes LLMs for role-playing interactions, incorporating self-confidence and early-stopping mechanisms in multi-round discussions to improve document reasoning and classification precision for triage tasks. Tang et al. (2023) uses multiple LLM-based agents as domain experts to collaboratively analyze medical reports and reach consensus on medical diagnoses. To the best of our knowledge, this is the first study to leverage LLM-MA for cohort selection in clinical trials. Our framework leverages LLM-MA agency to dynamically augment patient-trial matching with tailored knowledge, thereby enriching the contextual understanding of patient profiles and achieving better patient matching results.
3 Methods
We automate decomposed patient-matching workflows by utilizing the ability of LLMs to process and generate complex natural language. The proposed framework, illustrated in Figure 1, comprises five interconnected components, each with a distinct role in ensuring robust automation.
Knowledge Probing Agent: This component identifies knowledge gaps, both in terms of missing information in criteria and deficiencies in the LLM’s understanding of clinical trials and patient characteristics. By employing prompt-based self-probing techniques, inspired by SELFREF Kadavath et al. (2022) and MOREINFO Feng et al. (2023), the agent evaluates its own responses for correctness and consistency.If the agent determines that the original criterion is detailed and understandable enough for other agents, the criterion will be directly forwarded to the Matching Agent. Otherwise, the criterion is forwarded to the Navigation Agent for further processing.
Navigation Agent: In the MAKA framework, multiple specialized augmentation agents are available to enrich the criterion with domain-specific knowledge. The Navigation Agent determines the most suitable augmentation agent for the task and routes the workflow accordingly.
Knowledge Augmentation Agent: This component enriches the criterion with relevant, domain-specific information from different source. For instance, the Retrieval Agent Lewis et al. (2020) retrieves information from indexed medical databases, the Online Search Agent gathers supplementary knowledge from online sources, and the Self-Augment Agent leverages the LLM’s intrinsic knowledge to refine the original criterion.
Supervision Agent: Acting as a supervisor, this agent ensures that the augmented criterion aligns with the original instructions. For example, the augmented output must retain verbatim elements from the original criterion. This supervisory mechanism enhances the framework’s instruction-following capability and maintains consistency.
Matching Agent: Finally, with either the augmented or unaugmented criterion, as determined by the Knowledge Probing Agent, the Matching Agent evaluates patient eligibility in a zero-shot manner.
4 Evaluation
@@
| Zero-shot | CoT | MAKA | Zero-shot | CoT | MAKA | Zero-shot | CoT | MAKA | Zero-shot | CoT | MAKA | |
@@
| Average | 0.884 | 0.883 | 0.909 | 0.727 | 0.715 | 0.822 | 0.894 | 0.870 | 0.846 | 0.785 | 0.769 | 0.828 |
4.1 Data Description
@@
We use the dataset from Track 1 of the 2018 n2c2 challenge Stubbs et al. (2019), which is commonly used for evaluating patient matching in clinical trials. The dataset consists of 288 de-identified patients, each having between 2 to 5 unstructured longitudinal clinical notes in American English. The entire corpus contains 781,006 tokens, with an average of 2,711 tokens per patient. The challenge simulates a synthetic clinical trial with 13 predefined inclusion criteria, resulting in 3744 patient-criteria pairs in total. The detailed criteria can be found in Table 4. For each patient-criteria pair, a binary label is provided, indicating whether the patient "MET" or "NOT MET" the criterion. Two annotators with medical background labeled these data. The disagreements between the two annotators were resolved by an independent MD to establish the final ground truth labels.
4.2 Baseline and Metrics
@@
Previous research has demonstrated the superiority of LLM-based methods over traditional rule-based systems; hence, this study focuses exclusively on LLM-based benchmarks. We include two approaches: the zero-shot method with hand-crafted criteria Wornow et al. (2024), as shown in Table 5, and a Chain-of-Thought (CoT) approach Beattie et al. (2024) using the original criteria. In line with prior literature, we report accuracy, precision, recall, and F1-score for this binary classification task.
@@
Previous studies have primarily focused on evaluating classification performance at the individual criterion level, often overlooking the overall patient-trial assessment. In this study, we address this gap by examining patient-trial selection performance. Specifically, a patient is eligible for the trial only if all inclusion criteria are met. However, no patient in the n2c2 dataset satisfied all 13 criteria. To mitigate this, we synthesize a trail with selected criteria that had at least 100 patients meeting them, resulting in 28 eligible patients from a total cohort of 288. We then evaluated our method’s on this new synthetic trail in identifying eligible patients and reported the findings.
4.3 Results
4.3.1 Criterion-Level Performance
Table 1 shows the performance of three approaches—Zero-shot,CoT, MAKA—across multiple individual criteria. Our results indicate that while the CoT method slightly outperforms Zero-shot in Accuracy and Recall, MAKA consistently achieves higher scores, particularly in Accuracy (0.909) and Precision(0.822) on average. The F1-score of MAKA (0.828) suggests a balanced performance across criteria, although it falls slightly behind in a few specific cases such as "ABDOMINAL" and "English".
The MAKA-augmented criteria, presented in Table 6, demonstrate significant improvements over the baseline human-redefined criteria outlined in Table 5 from Wornow et al. (2024). These enhancements primarily focus on providing explicit clarifications and detailed expansions of key concepts, enabling LLMs to better align with trial requirements. For example, MAKA supplements normal ranges for criteria like "HBA1C" and "CREATININE" across diverse demographic groups. And they are also the criteria that benefit the most from MAKA’s augmentation. Additionally, MAKA provides a more comprehensive definition for criteria like "MAJOR-DIABETES," and use more widely recognized terminology, "heart attack", to explain clinical terms "Myocardial Infarction (MI)". These refinements enhance the ability of LLMs to match patients from clinical notes written in a variety of styles, resulting in notable performance gains for these criteria.
| Metric | Zero-shot | CoT | MAKA |

### v2

# Enhancing Clinical Trial Patient Matching through Knowledge Augmentation and Reasoning with Multi-Agents

@@
Abstract.
Matching patients effectively and efficiently for clinical trials is a significant challenge due to the complexity and variability of patient profiles and trial criteria. This paper introduces Multi-Agents for Knowledge Augmentation and Reasoning (MAKAR), a novel framework that enhances patient-trial matching by integrating domain-specific knowledge with structured reasoning. MAKAR consists of two key modules: the Augmentation Module and the Reasoning Module. The Augmentation Module enriches trial criteria by incorporating detailed explanations of relevant concepts to ensure clarity and completeness. The Reasoning Module then evaluates each health condition, following a structured, step-wise approach to determine eligibility and make the final matching decision. This paper enhances patient-trial matching by leveraging the agency and reasoning capabilities of large language models (LLMs) through automated agent interactions, including collaboration, critique, and navigation. Experimental results on a public dataset demonstrate that our framework surpasses existing benchmarks, achieving up to an 8% improvement in accuracy for specific criteria. Furthermore, in a real-world offline test, MAKAR achieved a 100% accuracy. These findings show MAKAR’s potential as a scalable and robust solution for clinical trial patient matching.
1. Introduction
Patient matching for clinical trials refers to the process of identifying and enrolling participants whose health profiles meet the specific eligibility criteria of a given clinical study. Efficient patient matching is crucial to accelerating the drug development process and reducing the overall cost of clinical trials. It also provides an opportunity for participants to access experimental treatments that could potentially improve their health outcomes and, in some cases, transform their quality of life.
@@ 1. Introduction
The emergence of of LLMs offers a promising opportunity to improve patient matching. LLMs, such as GPT-4 (Achiam et al., 2023), usually possess a large text window and advanced natural language understanding capabilities (Karanikolas et al., 2023), which enable them process complex trial criteria and patient data effectively. However, their general knowledge may not always be sufficient for knowledge-intensive tasks like patient matching in clinical trials, where domain-specific expertise is required. The knowledge gap inherent in LLMs and misssing information in the pre-defined criteria can limit their performance, highlighting the need for LLM augmentation of specialized medical knowledge (Shuster et al., 2021). While research indicates that LLMs do encode medical knowledge in a parameterized form (Singhal et al., 2023), this information can be outdated or misaligned with the special needs of certain clinical trials.
To address these limitations, we present a new framework called: Multi-Agents for Knowledge Augmentation and Reasoning (MAKAR). MAKAR is a novel framework designed to enhance the patient matching process. The framework consists of two core modules: the Augmentation Module and the Reasoning Module. The Augmentation Module enhances trial criteria by providing detailed explanations and includes three key components: a Router Agent, multiple Augmentation Agents, and a Critic Agent. The Reasoning Module systematically evaluates each criterion-related condition in a step-by-step manner to determine eligibility and make the final matching decision. It comprises a Reasoning Agent and a Supervision Matching Agent. By leveraging this multi-agent framework, MAKAR improves the performance on patient-trial matching by dynamically incorporating domain-specific knowledge and step-wise reasoning. Our experimental results demonstrate the effectiveness of MAKAR. On a public dataset, our framework achieves an average improvement of 3% over benchmark methods, with up to an 8% improvement for complex criteria. Furthermore, in an offline real-world data test, MAKAR outperformed the state-of-the-art closed-source reasoning LLMs and achieved 100% accuracy, showing its reliability in practical applications. Our key contributions are as follows:
We propose an automated method to refine and enhance trial criteria, making them more interpretable and suitable for AI-driven patient-trial matching.
We introduce a multi-agent framework that combines augmentation and reasoning, achieving strong performance across both public and real-world datasets.
We identify a gap in existing approaches and highlight the importance of explicitly incorporating well-structured, detailed explanations of related concepts to improve LLM-driven clinical reasoning.
2. Related Work
2.1. Clinical Trial Patient Matching
@@
Traditional patient matching systems rely on query-based structured databases, such as Criteria2Query (Yuan et al., 2019), or handcrafted rule-based approaches like RBC (Oleynik et al., 2019). However, these methods struggle to handle unstructured free-text data and are highly sensitive to variations in formatting and verbal expressions. Unfortunately, a significant portion of EHRs consists of free-text, including clinical notes, surgical records, radiology reports, pathology reports and so on (Sun et al., 2018). Thus, the effectiveness and scalability of traditional methods are significantly limited, making them unsuitable for future patient matching tasks.
@@
Recent advancements in LLMs have shown remarkable capabilities in directly processing free-text clinical trial criteria, offering scalable solutions without the need for extensive rule-based preprocessing. Many Studies have applied these models to patient-centric trial matching tasks using real-world data. For instance, Nievas et al. (2024) fine-tuned the open-source Llama 2 (Touvron et al., 2023) to automate patient matching. However, fine-tuning LLMs is computationally expensive, prompting alternative approaches that leverage frozen models. Jin et al. (2023) introduced TrialGPT, a patient-centric matching framework that assesses patient eligibility on a criterion-by-criterion basis, filtering out ineligible trials and ranking potentially eligible ones. Their results indicated a strong correlation between TrialGPT’s predictions and expert clinician judgments, achieving accuracy at the criterion level similar to human experts. Similarly, Gupta et al. (2024) proposed OncoLLM, an end-to-end pipeline that converts trial criteria into conversational prompts and uses a Chain-of-Thought (CoT) (Wei et al., 2022) reasoning strategy. OncoLLM outperforms GPT-3.5 while underperforms GPT-4, despite its smaller model size.
@@
In contrast, other studies have focused on trial-centric patient matching, typically using public deidentified datasets due to HIPAA-related data constraints. Wornow et al. (2024) developed a LLM-based system with a two-stage retrieval pipeline to identify suitable patients for trials using a zero-shot prompt approach. It reduces token processing costs while outperforming existing state-of-the-art models on the n2c2 2018 cohort selection benchmark (Stubbs et al., 2019). Likewise, Beattie et al. (2024) applied a CoT strategy in a Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) system using GPT-3.5 Turbo and GPT-4, demonstrating results comparable to those of Wornow et al. (2024).
@@
These works collectively highlight the potential of LLMs in clinical trial matching. However, existing approaches directly employ LLMs without verifying whether the trial criteria are sufficiently understood by the model (Webson and Pavlick, 2021) or whether there are gaps in the model’s knowledge (Lewis et al., 2020). Our work fits within the trial-centric patient matching literature and makes a novel contribution by automatically identifying and addressing these gaps, thereby enhancing the accuracy patient matching.
2.2. LLM-based Multi-Agents (LLM-MA)
@@
LLM-MA have seen rapid development across various domains, driven by the promising potential of LLMs as autonomous agents. Guo et al. (2024) provide a comprehensive overview of recent advancements in LLM-MA applications, categorizing them into two primary areas: Problem-Solving and World Simulation. Problem-solving LLM-MA systems explore the capabilities of agents with specialized expertise, enabling them to outperform single-agent LLM systems through collaborative efforts. For instance, Wang et al. (2024) proposed Mixture-of-Agents, a layered multi-agent architecture that achieves state-of-the-art performance in benchmarks like AlpacaEval 2.0 (Dubois et al., 2024).
In the healthcare domain, the application of LLM-MA systems is rapidly gaining traction. Yue et al.(Yue and Fu, 2024) introduce ClinicalAgent, a multi-agent architecture powered by LLM reasoning for multi-agent collaboration, achieving higher precision in predicting clinical trial outcomes. Similarly, Liao et al.(Liao et al., 2024) propose REFLECTool, a multi-agent framework that enables agents to learn how to select domain-specific tools, thereby extend their capabilities. TriageAgent(Lu et al., 2024) leverages LLMs for role-playing interactions, integrating self-confidence assessments and early-stopping mechanisms in multi-round discussions to improve document reasoning and classification accuracy for triage tasks. Tang et al.(Tang et al., 2023) use multiple LLM-based agents as domain experts to collaboratively analyze medical reports. And Li et al. (Li et al., 2024) further introduce the multi-agent collaboration framework to handle multimodal tasks.
To the best of our knowledge, this study is the first to apply an LLM-MA framework for cohort selection in clinical trials. Our approach explore the agency and reasoning capabilities of LLM-MA to dynamically enhance patient-trial matching with tailored knowledge. Our approach allows a better contextual understanding of patient profiles and lead to improved matching performance and more informed decision-making in clinical trial recruitment.
@@
| Zero-shot | CoT | MAKAR | Zero-shot | CoT | MAKAR | Zero-shot | CoT | MAKAR | Zero-shot | CoT | MAKAR | |
@@
| Average | 0.884 | 0.884 | 0.908 | 0.750 | 0.740 | 0.841 | 0.899 | 0.876 | 0.853 | 0.801 | 0.786 | 0.841 |
3. Methods
The proposed framework, shown in Figure 1, consists of two core modules: the Augmentation Module and the Reasoning Module. The augmentation module is designed to generate criteria with comprehensive explanations of all relevant concepts, while the reasoning module evaluates each condition within the criteria to determine eligibility and make the final matching decision. Additionally, the framework contains five interconnected components, each outlined below.
Router Agent In the MAKAR framework, multiple specialized augmentation agents are available to augment the criterion with domain-specific knowledge. The Router Agent determines the most suitable augmentation agent for the task and routes the workflow accordingly. Inspired by SELFREF (Kadavath et al., 2022) and MOREINFO (Feng et al., 2023), we adopt a prompt-based self-probing strategy. If the agent determines that the original criterion is detailed and understandable enough for other agents, the criterion will be directly forwarded to the Reasoning Module. Otherwise, the criterion is forwarded to the Knowledge Augmentation Agent chosen by the Router Agent for further processing.
Knowledge Augmentation Agent: This component augments the criterion with relevant, domain-specific information from different source. For instance, the Retrieval Agent (Lewis et al., 2020) retrieves information from indexed medical databases, the Online Search Agent gathers supplementary knowledge from online sources, and the Self-Augment Agent leverages the LLM’s intrinsic knowledge to refine the original criterion.
Critic Agent: Acting as a supervisor, this agent ensures that the augmented criterion is aligned with the original instructions. For instance, the original criterion must be preserved verbatim. Additionally, the critic agent verifies that the augmented output is well-structured (e.g., concepts are clearly itemized) to facilitate further step-wise reasoning. This critic mechanism are designed to improve the framework’s ability to follow instructions accurately while maintaining consistency and clarity.
Reasoning Agent With the augmented criterion and patient EHR data, the Reasoning Agent evaluates patient eligibility by assessing for each concepts mentioned in the augmented criterion. It generates a detailed reasoning summary, including assessments for all relevant health conditions, and forwards it to the Supervision Matching Agent.
Matching Agent: The Matching Agent determines patient eligibility by reviewing the reasoning assessment alongside the patient’s EHR data, making a final judgment in a zero-shot manner.
4. Evaluation & Results
Due to the sensitive nature of patient data, which is strictly protected under The Health Insurance Portability and Accountability Act (HIPAA), and the compliance requirements for processing detail clinical trial documents, full-scale online deployment is not currently feasible—particularly due to concerns about data security when using closed-source LLMs. Instead, to evaluate the effectiveness of our methods, we conducted two experiments: (1) an assessment using a public dataset and (2) a limited-scale offline evaluation with deidentified real-world data.
4.1. Data
@@
The public dataset we use is from Track 1 of the 2018 n2c2 challenge (Stubbs et al., 2019), which is commonly used for evaluating patient matching in clinical trials. The dataset consists of 288 de-identified patients, each having between 2 to 5 unstructured longitudinal clinical notes in American English. The entire corpus contains 781,006 tokens, with an average of 2,711 tokens per patient. The challenge simulates a synthetic clinical trial with 13 predefined inclusion criteria, resulting in 3744 patient-criteria pairs in total. The detailed criteria can be found in Table 6. For each patient-criteria pair, a binary label is provided, indicating whether the patient ”MET” or ”NOT MET” the criterion. Two annotators with medical background labeled these data. The disagreements between the two annotators were resolved by an independent MD to establish the final ground truth labels.
The criteria in the n2c2 dataset are relatively straightforward and the patient records are well-structured, making them easy to follow. To evaluate the potential of our proposed method in a more complex scenario, we constructed a real-world dataset. Specifically, we sampled 30 eligibility criteria—15 inclusion and 15 exclusion—from ongoing breast cancer trials listed on ClinicalTrials.gov. We then manually selected 10 patients who have been diagnosed with breast cancer from our in-house EHR system. For each patient, we extracted their most recent deidentified clinical visit record. An expert subsequently asseessed whether each patient met each criterion, resulting in 300 labeled criterion-patient pairs. Notably, if a patient’s information was insufficient to determine eligibility for a given criterion, they were labeled as ”NOT MET” to ensure a conservative patient selection. We listed the sampled criterion in Table 9
4.2. Baselines and Metrics
@@
Previous research has demonstrated the superiority of LLM-based methods over traditional rule-based systems; hence, this study focuses exclusively on LLM-based benchmarks. We include two approaches in Experiment 1: the zero-shot method with hand-crafted criteria (Wornow et al., 2024), as shown in Table LABEL:table:redefined_criteria, and a Chain-of-Thought (CoT) approach (Beattie et al., 2024) using the original criteria. In addtion, we also include o3-mini from OpenAI as the benchmark in Experiment 2 since reasoning LLMs have shown significant improvement on complex clinical scenarios(Xie et al., 2024). In line with prior literature, we report accuracy, precision, recall, and F1-score for this binary classification task.
@@
Previous studies using n2c2 dataset have primarily focused on evaluating classification performance at the individual criterion level, often overlooking the overall patient-trial assessment. In this study, we also examine patient-trial selection performance in Experiment 1. Specifically, a patient is eligible for the trial only if all inclusion criteria are met. However, no patient in the n2c2 dataset satisfied all 13 criteria. To mitigate this, we synthesize a trial with selected criteria that had at least 100 patients meeting them, resulting in 28 eligible patients from a total cohort of 288. We then evaluated our method’s on this new synthetic trial in identifying eligible patients and reported the findings.
4.3. Experiment Settings
Our experiments were conducted on an Amazon EC2 g5.16xlarge instance, equipped with a A10G GPU with 24 GB of memory. Unless otherwise specified, all agents in the framework are powered by GPT-4o.
For external knowledge augmentation, the Online Search Agent uses the Perplexity AI search engine to fetch up-to-date information from the web. The Retrieval Agent uses all-MiniLM-L12-v2, a pre-trained checkpoint from Sentence-Transformer (Reimers, 2019), as its encoder to construct the vector database of MedDRA and FDALabel.
| Metric | Zero-shot | CoT | MAKAR |